# Johannes Huemer [^2] These authors contributed equally to this work. Markus Murschitz33footnotemark: 3 Matthias Schörghuber Lukas Reisinger Thomas Kadiofsky Christoph Weidinger Mario Niedermeyer Benedikt Widy Marcel Zeilinger Csaba Beleznai Tobias Glück Andreas Kugi Patrik Zips44footnotemark: 4

### v1

@@ Abstract
Efficient material logistics play a critical role in controlling costs and schedules in the construction industry. However, manual material handling remains prone to inefficiencies, delays, and safety risks. Autonomous forklifts offer a promising solution to streamline on-site logistics, reducing reliance on human operators and mitigating labor shortages. This paper presents the development and evaluation of the Autonomous Dynamic All-terrain Pallet Transporter (ADAPT), a fully autonomous off-road forklift designed for construction environments. Unlike structured warehouse settings, construction sites pose significant challenges, including dynamic obstacles, unstructured terrain, and varying weather conditions. To address these challenges, our system integrates AI-driven perception techniques with traditional approaches for decision making, planning, and control, enabling reliable operation in complex environments. We validate the system through extensive real-world testing, comparing its long-term performance against an experienced human operator across various weather conditions. We also provide a comprehensive analysis of challenges and key lessons learned, contributing to the advancement of autonomous heavy machinery. Our findings demonstrate that autonomous outdoor forklifts can operate near human-level performance, offering a viable path toward safer and more efficient construction logistics.
@@ 1 Introduction
Digitalization and automation has the potential to revolutionize construction logistics. While digitalization allows optimal planning and tracking of materials, automation of machinery ensures fluid and timely material flow operation on the construction site. A multitude of reliable and flexible autonomous machines may be utilized to service the site in pull, push as well as just in time deliveries. In particular, forklifts are easy to bring to construction sites, even in large numbers, and are flexible in operation, which makes them well suited for ground-based material flow, lifting the workload of restricted cranes. When seamlessly integrated into the construction process through IT services, forklifts can greatly improve site efficiency by minimizing downtime, streamlined material handling, and preventing both equipment idleness and worker delays [5]. With the increasing shortage of skilled workers, the rough and dangerous working conditions, and the required increase in efficiency and decrease of errors, automation of these machines seems mandatory in the near future.
Figure 1: The autonomous off-road forklift ADAPT (Autonomous Dynamic All-terrain Pallet Transporter) demonstrating precision pallet loading onto a truck during a field demonstration. Throughout the development phase, the vehicle was equipped with additional hardware for testing outdoor sensor systems.
@@
However, the dynamic and unstructured environments of construction sites and the variable weather conditions pose significant challenges for automation. Unlike the controlled settings of autonomous warehouses, construction sites require flexible route planning and operation to navigate unpredictable terrains and obstacles. Safety is another paramount aspect, as 6% of total construction costs are due to accidents and one third of all fatalities at construction sites are caused by material handling equipment [6]. Despite advances in automation, there is a notable scarcity of autonomous machines capable of long-term operation in outdoor environments. Current solutions often fail in terms of robustness and adaptability, particularly when it comes to tasks such as unloading and loading full pallets from trucks. This gap underscores the need for innovative systems that can seamlessly integrate into the demanding conditions of construction sites.
@@
This paper presents the autonomous off-road forklift ADAPT (Autonomous Dynamic All-terrain Pallet Transporter), depicted in Figure 1, addressing the critical need for enhanced material handling solutions in construction. The scenario under consideration is a typical construction site that is not designed for autonomous operation. Predefined structures, guiding systems, or designated high-precision cargo locations are not required. Using advanced task and motion planning, purely synthetically trained geometry-based pallet detection, collision avoidance, and a new factor-graph-based joint vehicle localization and pallet mapping approach, our system aims to provide a reliable and efficient alternative to manual operation in this environment. The vehicle localization and pallet mapping approach is a specialized variant of traditional Simultaneous Localization and Mapping (SLAM) methods, specifically designed for manipulation tasks. Unlike most SLAM techniques, which rely on natural features (mid-level) or raw sensor data (low-level), this approach operates at a higher level by utilizing object poses—specifically, pallet locations. The overall approach and technology are designed to ensure mandatory adaptability to dynamic construction site conditions while maintaining robust and reliable operation. This is achieved through a combination of novel AI-based approaches (e.g., pallet detection) and classical physics-driven methodologies for task management, motion planning, and control. The use of primarily cost-effective sensors ensures affordability, while extensive real-world testing in various weather conditions validates the system’s performance. Through a comparative analysis with an experienced human operator, we demonstrate the potential of our automated forklift to give a realistic estimate of the performance and viability of autonomous machines in the coming years.
@@ 1.1 Contribution
  • Development and validation of an autonomous forklift system capable of loading and unloading operations in unstructured outdoor environments, demonstrating near-human performance across varied weather conditions, particularly robust operation under low to medium rainfall.
@@
Unlike many publications that focus solely on successful outcomes and scenarios, we provide a comprehensive section highlighting problems and key lessons learned from the long-term development and testing of autonomous machines. We believe that this will aid the scientific and engineering community in accelerating advances in heavy machine automation and, in the long run, contribute to safer and more reliable systems.
@@ 1.2 Article structure
The remainder of this paper is organized as follows. Section 2 presents a review of recent advances in autonomous outdoor machinery and related research in robotic material handling. Section 3 provides an overview of the proposed system, including its hardware and software components. The perception system, responsible for environment sensing, obstacle detection, and localization, is described in Section 4. Section 5 details the planning and control strategies used for autonomous navigation and load handling. In Section 6, we evaluate the system’s performance compared to an expert operator through experiments conducted in real-world outdoor environments and provide development and testing insights. Finally, Section 7 concludes the paper and discusses potential directions for future work. An overview of the system components and the corresponding sections is sketched in Fig. 2.
@@ 2 Related Work
While automated manipulation and logistics have reached commercial readiness in structured indoor environments, automated machines in rough, outdoor environments remain primarily in the research phase. Nevertheless, recent years have produced several promising prototypes. The HEAP platform [7] combines navigation and manipulation capabilities in a legged excavator. This outstanding research platform has demonstrated various tasks including excavation, dry-wall construction with natural stones, and forestry operations using both classical control techniques and novel AI approaches. Similarly, the Harveri platform [8] employs comparable technologies to harvest trees in challenging environments.
@@
Over the past two decades, several research platforms for autonomous forklift solutions have been published. Earlier work [10] explored vision-based handling tasks for autonomous outdoor forklifts, particularly focusing on vision systems for transporting molten aluminum in the metal industry. Another study [11] presented a system configuration for autonomous forklift operation that incorporated vision, laser range finders, sonar, and other sensors. This research analyzed the kinematics of a spin-turn mechanism and established essential system equations for path following based on time-varying feedback control law. However, in stark contrast to our work, their evaluation merely demonstrated the basic functionality of path following without presenting comprehensive real-world evaluations of loading success.
To our knowledge, only two similar scientific publications [12, 13] focus on autonomous forklifts in outdoor operation. The first [12] primarily emphasized motion planning with an anytime approach and explored new interaction modalities with human workers, such as voice commands. This work used trained AI only for object reacquisition after manual annotation of a single frame. Although it presented the robustness of the path planning, it did not provide a robustness or performance analysis of the complete loading cycle. Similarly, the second study [13] concentrated on the successful unloading of a specific pallet type from a truck, without offering comparative analysis against the performance of the human operator.
@@ 2.1 Localization, Mapping, and Traversability
To enable a system to operate in unfamiliar environments, a common approach is to use variants of SLAM (Simultaneous Localization and Mapping) [15]. SLAM is a technique that simultaneously constructs a map of an unknown environment while tracking the agent’s location within that environment, and has been extensively studied and applied in robotics applications [16, 17]. For autonomous unmanned forklifts, the SLAM method presented in [18] demonstrates robust operation in real-world, large-scale, and unstructured warehouse environments by employing a stereo-camera-based approach. This method achieves efficiency and robustness through a combination of direct (pixel-based) and indirect (feature-based) SLAM, focusing primarily on mapping indoor environments in mostly static settings.
Environment mapping is even more crucial for autonomous truck-mounted forklifts operating in unfamiliar construction sites. It serves as the foundation for traversability analysis and downstream path planning, enabling safe and efficient navigation. While mapping robot environments has been extensively studied over past decades, significant challenges remain in achieving human-like scene understanding—essential for enabling robots to navigate unstructured environments safely. Simple 2D occupancy grids may suffice for structured indoor spaces, but complex terrains require richer representations. The work [19] introduced a pioneering approach to address memory and runtime efficiency challenges in 3D occupancy maps, which has since been enhanced by methods such as spatio-temporal voxel layers [20] and VDB [21]. Other popular representations for 3D mapping include point clouds [22] and implicit representations [23, 24, 25].
@@ 2.2 Pallet Pose Estimation
Automated pallet manipulation requires the estimation of the 3D pallet pose with respect to the vision sensor. As we focus on a specific pallet type of known size and geometry (Euro-pallet), the vision task becomes instance-level pose estimation. To tackle this task, common pose-aware pallet detection approaches adopt either geometric cues from depth data or are based on visual appearance. Geometric schemes typically focus on 2D structure templates [34] or perform geometric fitting on point-cloud data from stereo [35], Time-of-Flight or LiDAR sensing [13]. In the presence of clutter and partial occlusions, however, geometric analysis schemes inherently become ambiguous. Appearance-based neural representations, sometimes complemented by depth in the form of RGB-D images, yield highly object-specific detectors, as demonstrated for 3D pose estimation in recent challenges [36]. Especially occlusion-robust local key-point representations [37, 38] offer accurate pose estimation results in high-resolution and semantically rich RGB or RGB-D data spaces. The work [39] introduces a pallet detection method based on the concept of Front Face Shots, which represent the access side of the pallets, where the pockets are. Their approach demonstrates that by combining a machine learning-based object detector with a kernel-based regression method, an accurate 6D pose can be computed, even for previously unseen pallet appearances.
@@ 2.3 Task Planning
Task Planning is essential for automating utility machinery, particularly in complex load manipulation tasks such as pallet handling. It involves high-level decision-making and sequencing of operations. Early approaches relied on finite state machines (FSMs) but evolved into hierarchical behavior trees for greater modularity and reactivity [41]. Similarly, the autonomous system under evaluation employs a modular behavior tree implementation to efficiently manage task execution. More recently, integrated Task and Motion Planning (TAMP) methods have emerged [42], combining symbolic reasoning with motion feasibility checks [43] or hierarchical architectures that leverage Linear Temporal Logic (LTL) for high-level planning and reactive behavior trees for low-level control [44]. Learning-based methods are being explored to infer planning domains from data, reducing the reliance on manually defined preconditions [45]. Recent developments have deployed LLMs for task planning. Although LLMs excel at processing natural language and common sense reasoning, directly translating abstract language inputs into executable plans is problematic due to their limited grounding in physical environments and inability to reason over complex task sequences [46]. LLMs struggle with the physical understanding of actions and fail to manage long-term dependencies in multiple steps, limiting their applicability to planning real-world tasks [47]. Recent research has explored the integration of LLMs with classical planners to capitalize on the language understanding of LLMs while leveraging the precision of PDDL-based planning [48, 49]. This hybrid approach improves the decision making capabilities of autonomous agents, allowing them to process natural language instructions and generate actionable plans using PDDL-based frameworks [50, 51]. However, no reliable long-term operation is yet feasible.
@@ 2.4 Motion Planning and Control
Various approaches have been developed to address these challenges. A Goal-Directed Rapidly exploring Random Tree (RRT) with multi-step refinement has been proposed for articulated construction machines [54]. Similarly, an extended Reeds-Sheep algorithm has been applied to wheel loader path planning [55]. A modified bug-like algorithm, integrated with Model Predictive Control (MPC) to ensure smoother trajectory execution, was proposed in [56]. Additionally, offline motion primitive generation combined with online receding-horizon planning has been successfully applied to tree harvester vehicles [57]. Application-focused path planning for yard automation was proposed in [58].
For tracking and control, a variant of the pure-pursuit method for the tracking of the path of articulated vehicles was introduced in [57], while another tracking controller has been specifically developed for articulated drum rollers that navigate construction sites [59].
@@
Beyond collision-free navigation, precise pallet docking is crucial for autonomous forklifts. Successful docking relies on accurate pallet pose estimation, achieved through vision-based systems, laser range finders, or hybrid sensor fusion approaches. The literature explores various motion planning techniques for this task, including geometric-based path generation [13, 61] and dynamic-based planning methods [62]. In addition, advanced control strategies, such as visual servoing [63], have been used to enhance the accuracy of jacking. Some methods also consider pallet inclination during insertion [64], improving adaptability and robustness in unstructured environments.
@@ 3.1 Vehicle Platform
Table 1 lists the main technical specifications of the BM154. The platform’s primary strengths lie in its advanced all-terrain capabilities, enabled by an all-wheel drive system and its center articulated steering with 3 rotational degrees of freedom, see Figure 3a. These features allow to navigate in uneven terrain under various ground conditions such as asphalt, gravel, mud, or grass. The platform consists of two chassis parts, the front part housing the motor and the hydraulic block, and the rear part incorporating the forks attached to the lifting mast. The motor side is designated as the main driving direction for autonomous operation, see Figure 4. This is motivated by considerations of load handling safety and operational visibility due to the unobscured view.
ADAPT has been extensively customized for automation since 2019 as part of several funded national and international research projects. Figure 3 illustrates the current hardware setup, highlighting the added components required for autonomous operation. The following sections provide a detailed overview of the hardware architecture and emphasize the differences between the automated prototype and the platform for manual operation.

### v2

@@ Abstract
Efficient material logistics play a critical role in controlling costs and schedules in the construction industry. However, manual material handling remains prone to inefficiencies, delays, and safety risks. Autonomous forklifts offer a promising solution to streamline on-site logistics, reducing reliance on human operators and mitigating labor shortages. This paper presents the development and evaluation of ADAPT (Autonomous Dynamic All-terrain Pallet Transporter), a fully autonomous off-road forklift designed for construction environments. Unlike structured warehouse settings, construction sites pose significant challenges, including dynamic obstacles, unstructured terrain, and varying weather conditions. To address these challenges, our system integrates AI-driven perception techniques with traditional approaches for decision making, planning, and control, enabling reliable operation in complex environments. We validate the system through extensive real-world testing, comparing its continuous performance against an experienced human operator across various weather conditions. Our findings demonstrate that autonomous outdoor forklifts can operate near human-level performance, offering a viable path toward safer and more efficient construction logistics.
@@ 1 Introduction
Digitalization and automation have the potential to revolutionize construction logistics. While digitalization allows optimal planning and tracking of materials, machinery automation ensures fluid and timely material flow operation on the construction site. A multitude of reliable and flexible autonomous machines may be utilized to service the site in pull, push, and just-in-time deliveries. In particular, forklifts are easy to bring to construction sites. They are flexible in operation, which makes them well suited for ground-based material flow, lifting the workload of restricted cranes. When seamlessly integrated into the construction process through IT services, forklifts can greatly improve site efficiency by minimizing downtime, streamlined material handling, and preventing both equipment idleness and worker delays [5]. With the increasing shortage of skilled workers, the rough and dangerous working conditions, and the required increase in efficiency and decrease in errors, automation of these machines seems mandatory.
Figure 1: The autonomous off-road forklift ADAPT (Autonomous Dynamic All-terrain Pallet Transporter) in its designated working environment — an unstructured construction site.
Image credit: © AIT/tm-photography.
@@
However, the dynamic and unstructured environments of construction sites and the variable weather conditions pose significant challenges for automation. Unlike the controlled settings of autonomous warehouses, construction sites require flexible route planning and operation to navigate unpredictable terrains and obstacles. Safety is another paramount aspect, as 6% of total construction costs are due to accidents, and one-third of all fatalities at construction sites are caused by material handling equipment [6]. Despite advances in automation, there is a notable scarcity of autonomous machines capable of long-term operation in outdoor environments. Current solutions often fail in robustness and adaptability, particularly regarding tasks such as unloading and loading full pallets from trucks. This gap underscores the need for innovative systems that can seamlessly integrate into the demanding conditions of construction sites.
@@
This paper presents the autonomous off-road forklift ADAPT (Autonomous Dynamic All-terrain Pallet Transporter), depicted in Figure 1. The scenario under consideration involves a truck-mounted forklift arriving on a typical construction site with a cargo truck. That site is not designed for autonomous operation; predefined structures, guiding systems, or designated high-precision cargo locations are not required. Using advanced task and motion planning, purely synthetically trained geometry-based pallet detection, collision avoidance, and a new factor-graph-based joint vehicle localization and pallet mapping approach, our system aims to provide a reliable and efficient alternative to manual operation in this environment. The vehicle localization and pallet mapping approach is a specialized variant of traditional Simultaneous Localization and Mapping (SLAM) methods specifically designed for manipulation tasks. Unlike most SLAM techniques, which rely on natural features (mid-level) or raw sensor data (low-level), this approach operates at a higher level by utilizing object poses—specifically, pallet locations. The overall approach and technology are designed to ensure mandatory adaptability to dynamic construction site conditions while maintaining robust and reliable operation. This is achieved through a combination of novel AI-based approaches (e.g., pallet detection) and classical physics-driven methodologies for task management, motion planning, and control. Using primarily cost-effective sensors ensures affordability, while extensive real-world testing in various weather conditions validates the system’s performance. Through a comparative analysis with an experienced human operator, we demonstrate the potential of our autonomous forklift to give a realistic estimate of the performance and viability of autonomous machines in the coming years. Since 100% operational reliability is not yet achievable for flexible use on construction sites, the forklift has been designed with a seamless human-machine interaction concept. Human workers, typically present on-site, can intervene when necessary for operational or safety reasons. After such interventions, autonomous operation resumes smoothly without disruption.
@@ 1.1 Contribution
  • Development of an autonomous forklift system capable of loading and unloading operations in unstructured outdoor environments, demonstrating near-human performance across varied weather conditions, particularly robust operation under low to medium rainfall.
@@ 1.2 Article Structure
The remainder of this paper is organized as follows. Section 2 reviews recent advances in autonomous outdoor machinery and related research in robotic material handling. Section 3 provides an overview of the proposed system, including its hardware and software components. The perception system, responsible for environment sensing, obstacle detection, and localization, is described in Section 4. Section 5 details the planning and control strategies for autonomous navigation and load handling. In Section 6, we evaluate the system’s performance compared to an expert operator through experiments conducted in real-world outdoor environments and provide development and testing insights. Finally, Section 7 concludes the paper and discusses potential directions for future work. An overview of the system components and the corresponding sections is sketched in Fig. 2.
@@ 2 Related Work
While automated manipulation and logistics are commercially viable in structured indoor environments, the deployment of autonomous machines in unstructured outdoor settings remains an active area of research. Nevertheless, recent years have produced several promising prototypes. The HEAP platform [7] combines navigation and manipulation capabilities in a legged excavator. This outstanding research platform has demonstrated various tasks including excavation, dry-wall construction with natural stones, and forestry operations using both classical control techniques and novel AI approaches. Similarly, the Harveri platform [8] employs comparable technologies to harvest trees in challenging environments.
@@
Over the past two decades, several research platforms for autonomous forklift solutions have been published. Earlier work [10] explored vision-based handling tasks for autonomous outdoor forklifts, particularly focusing on vision systems for transporting molten aluminum in the metal industry. Another study [11] presented a system configuration for autonomous forklift operation that incorporated vision, laser range finders, sonar, and other sensors. This research analyzed the kinematics of a spin-turn mechanism and established essential system equations for path-following based on time-varying feedback control law. However, in contrast to our work, their evaluation merely demonstrated the basic functionality of path following without presenting comprehensive real-world evaluations of loading success.
To our knowledge, only two similar scientific publications [12, 13] focus on autonomous forklifts in outdoor operation. The first [12] primarily emphasized motion planning with an anytime approach and explored new interaction modalities with human workers, such as voice commands. This work used trained AI only for object reacquisition after manual annotation of a single frame. Although it presented the robustness of the path planning, it did not provide a robustness or performance analysis of the complete loading cycle. Similarly, the second study [13] concentrated on successfully unloading a specific pallet type from a truck, without offering comparative analysis against the performance of the human operator.
@@ 2.1 Localization, Mapping, and Traversability
In order to enable a system to operate in unfamiliar environments, a common approach is to use variants of SLAM [15]. SLAM is a technique that simultaneously constructs a map of an unknown environment while tracking the agent’s location within that environment, and has been extensively studied and applied in robotics applications [16, 17]. For autonomous unmanned forklifts, the SLAM method presented in [18] demonstrates robust operation in real-world, large-scale, and unstructured warehouse environments by employing stereo-camera information. This method achieves efficiency and robustness through a combination of direct (pixel-based) and indirect (feature-based) SLAM, focusing primarily on mapping indoor environments in mostly static settings.
Environment mapping is even more crucial for autonomous truck-mounted forklifts operating in unfamiliar construction sites. It is the foundation for traversability analysis and downstream path planning, enabling safe and efficient navigation. While mapping robot environments has been extensively studied over past decades, significant challenges remain in achieving human-like scene understanding—essential for enabling robots to navigate unstructured environments safely. Simple 2D occupancy grids may suffice for structured indoor spaces, but complex terrains require richer representations. The work [19] introduced a pioneering approach to address memory and runtime efficiency challenges in 3D occupancy maps, which has since been enhanced by methods such as spatio-temporal voxel layers [20] and VDB (Volumetric Data Blocks) [21]. Other popular representations for 3D mapping include point clouds [22] and implicit representations [23, 24, 25].
@@ 2.2 Pallet Pose Estimation
Automated pallet manipulation requires estimating the 3D pallet pose with respect to the vision sensor. As we focus on a specific pallet type of known size and geometry (Euro-pallet), the vision task becomes instance-level pose estimation. Common pose-aware pallet detection approaches adopt geometric cues from depth data or rely on visual appearance. Geometric schemes typically focus on 2D structure templates [34] or perform geometric fitting on point-cloud data from stereo [35], Time-of-Flight or LiDAR sensing [13]. However, geometric analysis schemes inherently become ambiguous in the presence of clutter and partial occlusions. Appearance-based neural representations, sometimes complemented by depth in the form of RGB-D images, yield highly object-specific detectors, as demonstrated for 3D pose estimation in recent challenges [36]. Especially occlusion-robust local key-point representations [37, 38] offer accurate pose estimation results in high-resolution and semantically rich RGB or RGB-D data spaces. The work [39] introduces a pallet detection method based on the concept of Front Face Shots, which represent the accessible side of pallets with the fork pockets. Their approach demonstrates that by combining a machine learning-based object detector with a kernel-based regression method, an accurate 6D pose can be computed, even for previously unseen pallet appearances.
@@ 2.3 Task Planning
Task planning is essential for automating utility machinery, particularly in complex load manipulation tasks such as pallet handling. It involves high-level decision-making and sequencing of operations. Early approaches relied on finite state machines (FSMs) but evolved into hierarchical behavior trees for greater modularity and reactivity [41]. Similarly, the autonomous system under evaluation employs a modular behavior tree implementation to efficiently manage task execution. More recently, integrated Task and Motion Planning (TAMP) methods have emerged [42], combining symbolic reasoning with motion feasibility checks [43] or hierarchical architectures that leverage Linear Temporal Logic (LTL) for high-level planning and reactive behavior trees for low-level control [44]. Learning-based methods are explored to infer planning domains from data, reducing the reliance on manually defined preconditions [45]. Recent developments have deployed LLMs for task planning. Although LLMs excel at processing natural language and reasoning, directly translating abstract language inputs into executable plans is problematic due to their limited grounding in physical environments and inability to reason over complex task sequences [46]. LLMs struggle with the physical understanding of actions and fail to manage long-term dependencies in multiple steps, limiting their applicability to planning real-world tasks [47]. Recent research has explored the integration of LLMs with classical planners to capitalize on the language understanding of LLMs while leveraging the precision of PDDL-based planning [48, 49]. This hybrid approach improves the decision making capabilities of autonomous agents, allowing them to process natural language instructions and generate actionable plans using PDDL-based frameworks [50, 51]. However, no reliable long-term operation is yet feasible.
@@ 2.4 Motion Planning and Control
Various approaches have been developed to address these challenges. A Goal-Directed Rapidly exploring Random Tree (RRT) with multi-step refinement has been proposed for articulated construction machines [54]. Similarly, an extended Reeds-Sheep algorithm has been applied to wheel loader path planning [55]. A modified bug-like algorithm, integrated with Model Predictive Control (MPC) to ensure smoother trajectory execution, was proposed in [56]. Additionally, offline motion primitive generation and online receding-horizon planning have been successfully applied to tree harvester vehicles [57]. Application-focused path planning for yard automation was proposed in [58].
For tracking and control, a variant of the pure-pursuit method for the tracking of the path of articulated vehicles was introduced in [57]. Another tracking controller was specifically developed for articulated drum rollers that navigate construction sites [59].
@@
Beyond collision-free navigation, precise pallet docking is crucial for autonomous forklifts. Successful docking relies on accurate pallet pose estimation, achieved through vision-based systems, laser range finders, or hybrid sensor fusion approaches. The literature explores various motion planning techniques for this task, including geometric-based path generation [13, 61] and dynamic-based planning methods [62]. In addition, advanced control strategies, such as visual servoing [63], enhanced the accuracy of jacking. Some methods also consider pallet inclination during insertion [64], improving adaptability and robustness in unstructured environments.
@@ 3.1 Vehicle Platform
Table 1 lists the main technical specifications of the BM154. The platform’s primary strengths lie in its advanced all-terrain capabilities, enabled by an all-wheel drive system and its center-articulated steering with 3 rotational degrees of freedom, see Figure 3a. These features allow for navigating in uneven terrain under various ground conditions such as asphalt, gravel, mud, or grass. The platform consists of two chassis parts, the front part housing the motor and the hydraulic block, and the rear part incorporating the forks attached to the lifting mast. The motor side is designated as the main driving direction for autonomous operation, see Figure 4. This is motivated by considerations of load handling safety and operational visibility due to the unobscured view.
ADAPT has been extensively customized for automation since 2019 as part of several funded national and international research projects. Figure 3 illustrates the current hardware setup, highlighting the added components required for autonomous operation. The following sections provide a detailed overview of the hardware architecture and emphasize the differences between the autonomous prototype and the platform for manual operation.